A box that weighs 16N is resting on a horizontal surface. If a force of 6N is applied horizontally to the box and the coefficient of kinetic friction between the box and the surface is 0.4, the acceleration of the box will be:
Answer #1

- 6N EF friction (f) 16N = mg I IGN=mg x = 0.4 Net Force F-f=ma. as Frf a = 6- MR XIG 16/9.8 a=6-6.4x16 1.632 a = 6-6.4 = -0.